OpenStudy (czarluc):

how do you solve cos^2 x−cosx=0?

OpenStudy (anonymous):

\[\cos^2(x)-\cos(x)=0\] \[\cos^2(x)=\cos(x)\] *divide both sides by cos(x) \[\cos(x)=1\] turn 1 into cos with cos^-1 or arccos or shift cos, however you would like to call it. \[\cos(x)=\cos(0)\] You now can equate angles \[x=0+360k\] \[x=360k\]

OpenStudy (anonymous):

*About the division by cos(x) above, when you divide by a variable you must state it equals to zero, there lies your second answer: \[\cos(x)=0\] Same deal, turn 0 into cos: \[\cos(x)=\cos(90)\] Equate angles: \[x=90+180k\]

OpenStudy (anonymous):

The answers: \[x_1=360k\] \[x_2=90+180k\] The addition of 360k and 180k just takes into account that x has infinite solutions every 360 degrees

OpenStudy (anonymous):